OpenStudy (anonymous):

Evaluate each function rule for x = -6.

OpenStudy (anonymous):

7. y = 7x + 4

OpenStudy (nirmalnema):

put the value of x in y = 7x + 4 y = 7(-6) + 4 y=-42+4 =-38

OpenStudy (anonymous):

so the answer is y=-42 + 4 = -38 then? :) @nirmalnema